A truck heading east has an initial velocity of 6 m/s. It accelerates at 2 m/s2 for 12 seconds. What distance does the truck travel in the given time?
1
Expert's answer
2020-06-25T09:34:56-0400

The kinamatic law for the uniform acceleration is the following:


"d = v_0t + \dfrac{at^2 }{2}"

Here "d" is the distance traveled by the truck, "a" is its acceleration and "t" is time.

Substituting numerical values, get:


"d = 6m/s\cdot 12s+\dfrac{2m/s^2\cdot 12^2s^2}{2} = 216 m"

Answer. d = 216m